A car moving with a velocity of 54 km/h accelerates uniformly at the rate of 2 m/s2. Calculate the distance travelled from the place where the acceleration began to that where the velocity reaches 72 km/h, and the time taken to cover this distance.
1
Expert's answer
2020-10-12T11:30:25-0400

let's first convert km/hr into m/s

54 km/hr : 15m/s

72 km/hr : 20m/s

Applying :

"v^2-u^2=2aS"

"20\times20-15\times15=2\times2\times S\\S=175/4=43.75 m"